[Ncep.nhc.nco_contacts] TSB Morning Rounds - Wed Apr 26, 2017

TSB Admin nhc.tsbadmin at noaa.gov
Wed Apr 26 15:48:02 UTC 2017


Here is today's summary of NHC computer operations:
--- Matt Sardi

----------

TAFB:
1. STARFISH (Atlantic desk) froze up this morning and had to be rebooted.  
TSB will open a Helpdesk ticket if problems continue.

2. CUMULUS (Surface desk) and BALLYHOO (Atlantic desk) have been performing  
much better since TSB (Matt Sardi) identified typos in the xorg.conf  
configuration files on both boxes, which NCO has now fixed.
HSU:
1. WFO storm surge testing continues through Thursday. On each day, Mike  
Brennan submits a new test psurge run and the affected WFOs verify ingest.  
Testing has gone well so far.

Other items:
1. CARCAH/TSB continue to work with NCO to resolve configuration issues on  
the scout2 RHEL6 replacement workstation.

2. Due to today's CWD, NCO's upgrade on the core Internet routers in  
College Park, MD has been postponed until Thursday, April 27, 2017.

The intermittent impacts during the 1015 UTC (6:15 am ET) to 1120 UTC (7:20  
am ET) period will affect NHC and other Remote NCEP Centers:
· Loss of internet service
· Interruption to NCEP VPN access and WCOSS
· Delay of NCEP-hosted email services, such as listservs

There will be a second period of potential impact from 1900 UTC (3pm ET) to  
2015 UTC (4:15 pm ET). The work during this 75 minutes should not result in  
any outages, but it is a high risk period of work that they want users to  
be aware of.
NHC RFCs:
None.

-------------- next part --------------
An HTML attachment was scrubbed...
URL: https://www.lstsrv.ncep.noaa.gov/pipermail/ncep.nhc.nco_contacts/attachments/20170426/fa0306e6/attachment.html 


More information about the Ncep.nhc.nco_contacts mailing list